A 3.0 mH inductor is connected in parallel with a variable capacitor. The capacitor can be varied from 120 pF to 320 pF . a) What is the minimum oscillation frequency for this circuit? b) What is the maximum oscillation frequency for this circuit?
Answer #1

We know that oscillation frequency

f=1/(2π√(LC))

​​a)for minimum oscillation frequency

L=3.0mH =3×10-3 H, C=320×10-12 F

fmin=1/(2π√(3×10-3×320×10-12))

fmin=106/(2π√(0.96)

fmin=162436.83 Hz Answer

b)for maximum frequency C=120pF

fmax=1/(2π√(3×10-3×120×10-12)

fmax=106/(2π√0.36)

fmax=265258.24 Hz. Answer